OpenStudy (anonymous):

Solve n equals four times r for r. r=4 times n r equals four times r r equals n plus four r equals n divided by four

OpenStudy (anonymous):

B I think.

OpenStudy (anonymous):

r equals n divided by four because u divide both sides by 4to get r by itself

OpenStudy (anonymous):

n=4*r r=n/4